Suppose you have a 1020 steel plate in axial loading, with an axial F of 30000 lbf.  The plate is 1⁄4” thick, 2” high, and 6” length. The hole is in the center of the  plate, and is 1/4“ in diameter. The plate is at room temperature, and the strain at  fracture is greater than εf > 0.1. If Sy = 80 ksi and Sut = 100 ksi, calculate the highest  stress in the plate.


1
Expert's answer
2021-06-10T03:46:48-0400

There is no need for a stress concentration factor as this material is ductilenote that ε

f > 0.1

The highest stress will be at the lowest cross-sectional area, which occurs at

the hole. Therefore

"\sigma= \frac{
F}{
(t)(h - d)}
=
\frac{30000lbf}{
(0.25in)(2in - 0.25in)}
= 68,600psi = 68.6ksi"